What is the focal length of the lens in your eye when you read a book that is 35 cm from your eye? The distance from the lens to the retina is 19mm.
1
Expert's answer
2011-05-31T07:14:13-0400
Using thin lens formula we can find the focal length:
1/f = 1/S1 + 1/S2

f=(S2*S1)/(S2+S1)

f=(35*1.9)/36.9≈1.8 cm

Answer :1.8 cm